OpenStudy (anonymous):

ok when writing an equation of a line that passes through the points (0,4) and (4,0) my answer came out as 0/4 so isn't this undefined?

OpenStudy (anonymous):

0/4=0, but that is not the slope. The slope is (y2-y1)/(x2-x1)=(0-4)/(4-0)=-4/4=-1. Then we use point slope form to obtain y-4=-(x-0). so y=-x+4